OpenStudy (anonymous):

lim [1/(x+3)4]=infinity. Prove x→−3

OpenStudy (anonymous):

I don't understand limits

OpenStudy (anonymous):

CORRECTION. It should be lim as x approaches -3 [1/ (x+3)^4) = infinity

OpenStudy (anonymous):

is this \[\lim_{x\rightarrow -3}\frac{1}{(x+3)^4}\]?

OpenStudy (anonymous):

YES and then it equals infinity

OpenStudy (anonymous):

do you want to understand why it is true?

OpenStudy (anonymous):

yes. like my teacher writes some statement using epsilon and delta.

OpenStudy (anonymous):

then proves it.

OpenStudy (anonymous):

cause 1/0 is infinity

OpenStudy (anonymous):

ok before we worry about that, we know that x cannot be equal to -3 because you cannot divide by 0. so suppose x is say -2.9 then \[-2.9+3=.1\] and \[.1^4=.0001\] and finally \[\frac{1}{.0001}=10000\]

OpenStudy (anonymous):

just for the record \[\frac{1}{0}\] is not a number and it is not infinity either

OpenStudy (anonymous):

he doesnt jsut say that 1/0 is infinity. He will write liek for any M is greater than 0, we want delta greater tahn 0 such that then writes some statements with absolute values. At teh end, teh satement shud be like choose delta = ??

OpenStudy (anonymous):

in any case what we see is that as x gets close to -3, \[\frac{1}{(x+3)^4}\] will grow bigger and bigger

OpenStudy (anonymous):

I'm confused how to do teh same thing with this problem. I'm not sure if anyone else learned it this way.

OpenStudy (anonymous):

yes i understand what you want, i just wanted to give an example. in my example N was 10000 and delta was 0.1

OpenStudy (anonymous):

oh ok.

OpenStudy (anonymous):

suppose you want it to be bigger than say 1,000,000,000 what delta would you choose?

OpenStudy (anonymous):

delta = 10000

OpenStudy (anonymous):

hmmm i see we need some work here. delta is supposed to be the distance between x and -3 i certainly am not going to pick a very large number. it will be a very small number. like in my example i had N at 10000 and delta turned out to be only 0.1 so we expect a nice small delta. the bigger N, the smaller delta will be

OpenStudy (anonymous):

that is the larger the output, the closer x will have to be to -3

OpenStudy (anonymous):

lets solve it for N =1,000,000,000

OpenStudy (anonymous):

yeah i mean when i insert the -3 into the equation i find 1 div 0 that cant solve so we have to insert value approx -3 in order to get finit

OpenStudy (anonymous):

i want \[\frac{1}{(x+3)^4}>1,000,000,000\]

OpenStudy (anonymous):

to solve this inequality for x, i would take the reciprocal of both sides and write \[(x+3)^4<0.000 000001\]

OpenStudy (anonymous):

then take the 4th root of both sides to get \[-.01<x+3<.01\]

OpenStudy (anonymous):

in absolute value notation this says \[|x+3|<0.01\]

OpenStudy (anonymous):

good explanation satellite

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

so to get what you need, replace 1000000000 by N and the find the appropriate delta the steps are identical except with N instead of the number

OpenStudy (anonymous):

start with \[\frac{1}{(x+3)^4}>N\] then \[(x+3)^4<\frac{1}{N}\] then \[-\frac{1}{\sqrt[4]{N}}<x+3<\frac{1}{\sqrt[4]{N}}\] finally \[|x+3|<\frac{1}{\sqrt[4]{N}}\]and you are done

OpenStudy (anonymous):

So how do i write that for any N greater than 0, we want delta greater tahn 0 such taht ...then use absolute value statements?

OpenStudy (anonymous):

For ex. a prob we did was lim as x approaches 0, 1/x^2 = infinity. So the statement was for any M greater than 0, we want delta greater tahn 0 so taht absolute value of x is less than delta implies that 1/x^2 is greater tahn M.

OpenStudy (anonymous):

Do u understand what i mean??This is what i wrote but as an eqation lim 1/x^2 = infinity. x app. 0 For any \[M > \delta\] we want \[\delta > 0\] so that \[\left| x \right| < \delta \implies that 1/x ^{2} >M\] \[1/X ^{2} > M \implies x ^{2} < 1/M \implies \sqrt{x ^{2}} < \sqrt{1/M} \] \[\left| X \right| < 1/\sqrt{m} \] Choose \[\delta = 1/\sqrt{M}\]

OpenStudy (anonymous):

yes it is what i wrote above exactly, but i used N instead of M and found \[\delta =\frac{1}{\sqrt[4]{N}}\] would do

OpenStudy (anonymous):

btw for when you say the limit it is infinity you do not usually say "for N > 0" because you are concerned with large N. you usually say "for any N no matter how large" or something like that

OpenStudy (anonymous):

oh okay. Is this correct?? For any for any N no matter how large, we want delta greater tahn 0, so taht \[\left| x+3 \right| <\delta\] .....etc (what u wrote)